### v2

# Mean Estimation in Banach Spaces Under Infinite Variance and Martingale Dependence

Abstract
We consider estimating the shared mean of a sequence of heavy-tailed random variables taking values in a Banach space. In particular, we revisit and extend a simple truncation-based mean estimator first proposed by Catoni and Giulini. While existing truncation-based approaches require a bound on the raw (non-central) second moment of observations, our results hold under a bound on either the central or non-central $p$th moment for some $p\in(1,2]$. Our analysis thus handles distributions with infinite variance. The main contributions of the paper follow from exploiting connections between truncation-based mean estimation and the concentration of martingales in smooth Banach spaces. We prove two types of time-uniform bounds on the distance between the estimator and unknown mean: line-crossing inequalities, which can be optimized for a fixed sample size $n$, and iterated logarithm inequalities, which match the tightness of line-crossing inequalities at all points in time up to a doubly logarithmic factor in $n$. Our results do not depend on the dimension of the Banach space, hold under martingale dependence, and all constants in the inequalities are known and small.
1 Introduction
Mean estimation is perhaps the most important primitive in the statistician’s toolkit. When the data is light-tailed (perhaps sub-Gaussian, sub-Exponential, or sub-Gamma), the sample mean is the natural estimator of this unknown population mean. However, when the data fails to have finite moments, the naive plug-in mean estimate is known to be sub-optimal.
The failure of the plug-in mean has led to a rich literature focused on heavy-tailed mean estimation. In the univariate setting, statistics such as the thresholded/truncated mean estimator [36, 20], trimmed mean estimator [32, 27], median-of-means estimator [31, 21, 2], and the Catoni M-estimator [5, 38] have all been shown to exhibit favorable convergence guarantees. When a bound on the variance of the observations is known, many of these estimates enjoy sub-Gaussian rates of performance [24], and this rate gracefully decays when only a bound on the $p$th central moment is known for some $p>1$ [3].
In the more challenging setting of multivariate heavy-tailed data, modern methods include the geometric median-of-means estimator [30], the median-of-means tournament estimator [26], and the truncated mean estimator [6]. We provide a more detailed account in Section 1.2.
Of the aforementioned statistics, the truncated mean estimator is by far the simplest. This estimator, which involves truncating observations to lie within an appropriately-chosen ball centered at the origin, is extremely computationally efficient and can be updated online, very desirable for applied statistical tasks. However, this estimator also possesses a number of undesirable properties. First, it is not translation invariant, with bounds that depend on the raw moments of the random variables. Second, it requires a known bound on the $p$th moment of observations for some $p\geq 2$, thus requiring that the observations have finite variance. Third, bounds are only known in the setting of finite-dimensional Euclidean spaces — convergence is not understood in the setting of infinite-dimensional Hilbert spaces or Banach spaces.
The question we consider here is simple: are the aforementioned deficiencies fundamental to truncation-based estimators, or can they be resolved with an improved analysis? The goal of this work is to show that the latter is true, demonstrating how a truncation-based estimator can be improved to handle fewer than two central moments in general classes of Banach spaces.
1.1 Our Contributions
In this work, we revisit and extend a simple truncation-based mean estimator due to Catoni and Giulini [6]. Our estimator works by first using a small number of samples to produce a naive mean estimate, say through a sample mean. Then, the remaining sequence of observations is truncated to lie in an appropriately-sized ball centered at this initial mean estimate. Centering the remaining observations in this way is what enables bounds that depend on the centered $p$-th moment instead of the raw moment. The truncated samples are then averaged to provide our final estimate.
While existing works study truncation-based estimators via PAC-Bayesian analyses [6, 11, 22], we find it more fruitful to study these estimators using tools from the theory of Banach space-valued martingales. In particular, by proving a novel extension of classical results on the time-uniform concentration of bounded martingales due to Pinelis [33, 34], we are able to greatly improve the applicability of truncation-based estimators. In particular, our estimator and analysis improves over that in [6] in the following ways:
  1. 1.
    The analysis holds in arbitrary 2-smooth Banach spaces instead of just finite-dimensional Euclidean space. This not only includes Hilbert spaces but also the commonly-studied $L^{\alpha}$ and $\ell^{\alpha}$ spaces for $2\leq\alpha<\infty$.
  2. 2.
    Our results require only a known upper bound on the conditional central $p$th moment of observations for some $p>1$, and are therefore applicable to data lacking finite variance. Existing bounds for truncation estimators, on the other hand, require a bound on the non-central second moment.
  3. 3.
    Our bounds hold for data with a martingale dependence structure, not just for independent and identically distributed data.
  4. 4.
    Our bounds are time-uniform. We prove two types of inequalities: line-crossing inequalities, which can be optimized for a target sample size, and law of the iterated logarithm (LIL) inequalities, which match the tightness of the boundary-crossing inequalities at all times simultaneously up to a doubly logarithmic factor in sample size.
  5. 5.
    We show that our estimator exhibits strong practical performance, and that our derived bounds are tighter than existing results in terms of constants. We run simulations which demonstrate that, for appropriate truncation diameters, the distance between our estimator and the unknown mean is tightly concentrated around zero.
Informally, if we assume that the central $p$th moments of all observations are conditionally bounded by $v$, and we let $\widehat{\mu}_{n}$ denote our estimate after $n$ samples, then we show that
$\|\widehat{\mu}_{n}-\mu\|=O\left(\beta v^{1/p}(\log(1/\delta)/n)^{\frac{p-1}{p}}\right)\qquad\text{with probability }\geq 1-\delta,$
where $\beta$ is a parameter governing the smoothness of the Banach space. As far as we are aware, the only other estimator to obtain the same guarantee in a similar setting is Minsker’s geometric median-of-means [30]. (While he doesn’t state this result explicitly, it is easily derivable from his main bound—see Appendix B for the details.) Minsker also works in a Banach space, but assumes that it is separable and reflexive, whereas we will assume that it is separable and smooth; the latter appears stronger. While we obtain the same rates, we feel that our truncation-style estimator has several benefits over geometric median-of-means. Ours is computationally lightweight and easy to compute exactly. Importantly, we handle martingale dependence while Minsker does not, thus allowing our estimator to provide time-uniform guarantees. Finally, our analysis is significantly different from Minsker’s—and from existing analyses of other estimators under heavy-tails—and may be of independent interest.
1.2 Related Work
Section 1.1 discussed the relationship between this paper and the two most closely related works of Catoni and Giulini [6] and Minsker [30]. We now discuss how our work is related to the broader literature, none of which addresses our problem directly, but tackles simpler special cases of our problem (e.g., assuming more moments or boundedness, or with observations in Hilbert spaces or Euclidean spaces).
Heavy-tailed mean estimation under independent observations
Truncation-based (also called threshold-based) estimators have a rich history in the robust statistics literature, dating back to works from Tukey, Huber, and others [20, 36]. These estimators have either been applied in the univariate setting or in $\mathbb{R}^{d}$ as in Catoni and Giulini [6]. A related estimator is the so-called trimmed-mean estimator, which removes extreme observations and takes the empirical mean of the remaining points [32, 27]. For real-valued observations with finite variance, the trimmed-mean has sub-Gaussian performance [32].
Separately, Catoni and Giulini [5] introduce an approach for mean estimation in $\mathbb{R}^{d}$ based on M-estimators with a family of appropriate influence functions. This has come to be called “Catoni’s M-estimator.” It requires at least two moments and fails to obtain sub-Gaussian rates. It faces the the additional burden of being less computationally efficient. A series of followup works have improved this estimator in various ways: for real-valued observations, Chen et al. [7] extend it to handle a $p$-th moment for $p\in(1,2)$, Gupta et al. [15] refine the constants in the preceding paper, and Mathieu [29] studies the optimality of general M-estimators for mean estimation.
Another important line of work on heavy-tailed mean estimation is based on median-of-means estimators [31, 21, 2]. These estimators generally break a dataset into several folds, compute a mean estimate on each fold, and then compute some measure of central tendency amongst these estimates. For real-valued observations, Bubeck et al. [3] study a median-of-means estimator that holds under infinite variance. Their estimator obtains the same rate as ours and Minsker’s. Most relevant for our work is the result on geometric median-of-means due to Minsker [30], which can be used to aggregate several independent mean estimates in general separable Banach spaces. In Hilbert spaces, when instantiated with the empirical mean under a finite variance assumption, geometric median-of-means is nearly sub-Gaussian (see discussion in Section 1.1). We compare our threshold-based estimator extensively to geometric median-of-means in the sequel and demonstrate that we obtain the same rate of convergence.
[…]
While the original form of the tournament median-of-means estimator was computationally inefficient (with computation hypothesized to be NP-Hard in a survey by Lugosi and Mendelson [24]), a computationally efficient approximation was developed by Hopkins [16], with followup work improving the running time [8]. Tournament median-of-means was extended to general norms in $\mathbb{R}^{d}$ [25], though the authors note that this approach is still not computationally feasible. Median-of-means style approaches have also been extended to general metric spaces [19, 9]. Of the above methods, only the geometric median-of-means estimator can handle observations that lack finite variance.
Sequential concentration under martingale dependence
Time-uniform concentration bounds, or concentration inequalities that are valid at data-dependent stopping times, have been the focus of significant recent attention [17, 18, 40]. Such results are often obtained by identifying an underlying nonnegative supermartingale and then applying Ville’s inequality [37], a strategy that allows for martingale dependence quite naturally. This approach is also used here. Wang and Ramdas [38] extend Catoni’s M-estimator to handle both infinite variance and martingale dependence in $\mathbb{R}$, while Chugg et al. [11] give a sequential version of the truncation estimator in $\mathbb{R}^{d}$, though they require a central moment assumption and finite variance. The analyses of both Catoni and Giulini [6] and Chugg et al. [11] rely on so-called “PAC-Bayes” arguments [4, 12]. Intriguingly, while we analyze a similar estimator, our analysis avoids such techniques and is much closer in spirit to Pinelis-style arguments [33, 34].
Howard et al. [17, 18] provide a general collection of results on time-uniform concentration for scalar processes, which in particular imply time-uniform concentration results for some heavy-tailed settings (e.g. symmetric observations). Likewise, Whitehouse et al. [40] provide a similar set of results in $\mathbb{R}^{d}$. While interesting, we note that these results differ from our own in that they are self-normalized, or control the growth of a process appropriately normalized by some variance proxy (there a mixture of adapted and predictable covariance). The results also don’t apply when only a bound on the $p$th moment is known, and the latter set of results have explicit dependence on the ambient dimension $d$.
Concentration in Hilbert and Banach Spaces
There are several results related to concentration in infinite-dimensional spaces. A series of works has developed self-normalized, sub-Gaussian concentration bounds in Hilbert spaces  [39, 1, 10] based on the famed method of mixtures [13, 14]. These results have not been extended to more general tail conditions. Significant progress has been made on the concentration of bounded random variables in smooth and separable Banach spaces. Pinelis [33, 34] presented a martingale construction for bounded observations, thus enabling dimension-free Hoeffding and Bernstein inequalities. Dimension-dependence is replaced by the smoothness parameter of the Banach space, which for most practical applications (in Hilbert spaces, say) equals one. These results were strengthened slightly by Howard et al. [17]. Recently, Martinez-Taboada and Ramdas [28] gave an empirical-Bernstein inequality in Banach spaces, also using a Pinelis-like construction. Our work adds to this line by extending Pinelis’ tools to the heavy-tailed setting.
1.3 Preliminaries
We introduce some of the background and notation required to state our results. We are interested in estimating the shared, conditional mean $\mu$ of a sequence of random variables $(X_{n})_{n\geq 1}$ living in some separable Banach space $(\mathbb{B},\|\cdot\|)$. Recall that a Banach space is a complete normed vector space; examples include Hilbert spaces, $\ell^{\alpha}$ sequence spaces, and $L^{\alpha}$ spaces of functions. We make the following central assumption.
Assumption 1.
We assume $(X_{n})_{n\geq 1}$ are a sequence of $\mathbb{B}$-valued random variables adapted to a filtration $\mathcal{F}\equiv(\mathcal{F}_{n})_{n\geq 0}$ such that
  1. (1)
    $\mathbb{E}(X_{n}\mid\mathcal{F}_{n-1})=\mu$, for all $n\geq 1$ and some unknown $\mu\in\mathbb{B}$, and
  2. (2)
    $\sup_{n\geq 1}\mathbb{E}\left(\|X_{n}-\mu\|^{p}\mid\mathcal{F}_{n-1}\right)\leq v<\infty$ for some known constants $p\in(1,2]$ and $v>0$.
The martingale dependence in condition (1) above is weaker than the traditional i.i.d. assumption, requiring only a constant conditional mean. This is useful in applications such as multi-armed bandits, where we cannot assume that the next observation is independent of the past. Meanwhile, condition (2) allows for infinite variance, a weaker moment assumption than past works studying concentration of measure in Banach spaces (e.g., [30, 33, 34]). In Appendix C we replace condition (2) with a bound on the raw moment (that is, $\mathbb{E}(\|X_{n}\|^{p}|\mathcal{F}_{n-1})$) for easier comparison with previous work. We note that other works studying truncation-based estimators have exclusively considered the $p\geq 2$ setting where observations admit covariance matrices [11, 6, 24]. We focus on $p\in(1,2]$ in this work, but it is likely our techniques could be naturally extended to the $p\geq 2$ setting. We leave this as future work.
In order obtain concentration bounds, we must assume the Banach space is reasonably well-behaved. This involves assuming that is it both separable and smooth. A space is separable if it contains a countable, dense subset, and a real-valued function $f:\mathbb{B}\to\mathbb{R}$ is $(2,\beta)$-smooth if, for all $x,y\in\mathbb{B}$, $f(0)=0$, $|f(x+y)-f(x)|\leq\|y\|$, and
$f^{2}(x+y)+f^{2}(x-y)\leq 2f^{2}(x)+2\beta^{2}\|y\|^{2}.$ (1.1)
We call a Banach space as $\beta$-smooth if its norm is $(2,\beta)$-smooth. [^1] In the parlance of convex analysis, $(2,\beta)$-smoothness is equivalent to $f^{2}$ being $2\beta^{2}$-smooth, i.e. $f^{2}(tx+(1-t)y)\geq tf^{2}(x)+(1-t)f^{2}(y)-\beta^{2}t(1-t)\|x-y\|^{2}.$ Observe that $\beta\geq 1$ when $f$ is the norm, which can be seen by taking $x=0$ in (1.1).
Assumption 2.
We assume that the Banach space $(\mathbb{B},\|\cdot\|)$ is both separable and $\beta$-smooth.
Assumption 2 is common when studying concentration of measure in Banach spaces [33, 34, 17, 28]. The following result assures us that the sample mean does in fact concentrate in smooth Banach spaces, which will be required in our analysis of the naive mean estimate we use as a centering device (see $\widehat{Z}_{k}$ in Section 2). The proof in Appendix A relies on an argument based on decoupled tangent sequences, which may be of independent interest. If the underlying space $\mathbb{B}$ is a Hilbert spaces and $p=2$, the multiplicative factor of 2 in the bound below can be dropped.
Lemma 1.1 (Naive Mean Concentration).
Let $(X_{n})_{n\geq 1}$ be a sequence random variables satisfying Assumption 1 taking values in a Banach space satisfying Assumption 2. Then, for any $\delta\in(0,1)$ and $n\geq 1$, we have
$\mathbb{P}\left(\|\widehat{\mu}_{n}-\mu\|\leq\frac{2\beta v^{1/p}}{\delta^{1/p}n^{(p-1)/p}}\right)\geq 1-\delta,$
where $\widehat{\mu}_{n}:=n^{-1}\sum_{m=1}^{n}X_{m}$ denotes the usual sample mean.
We emphasize that $\beta$ is not akin to the dimension of the space. For instance, infinite-dimensional Hilbert spaces have $\beta=1$ and $L^{\alpha}$ and $\ell^{\alpha}$ spaces have $\beta=\sqrt{\alpha-1}$ for $\alpha\geq 2$. Thus, bounds which depend on $\beta$ are still dimension-free.
Notation and background
For notational simplicity, we define the conditional expectation operator $\mathbb{E}_{n-1}[\cdot]$ to be $\mathbb{E}_{n-1}[X]:=\mathbb{E}(X\mid\mathcal{F}_{n-1})$ for any $n\geq 1$. If $S\equiv(S_{n})_{n\geq 0}$ is some stochastic process, we denote the $n$-th increment as $\Delta S_{n}:=S_{n}-S_{n-1}$ for any $n\geq 1$. For any process or sequence $a\equiv(a_{n})_{n\geq 1}$, denote by $a^{n}$ the first $n$ values: $a^{n}=(a_{1},\dots,a_{n})$. We say the process $S$ is predictable with respect to filtration $\mathcal{F}$, if $S_{n}$ is $\mathcal{F}_{n-1}$-measurable for all $n\geq 1$. Our analysis will make use of both the Fréchet and Gateaux derivatives of functions in a Banach space. We do not define these notions here; see Ledoux and Talagrand [23].
Outline
Section 2 provides statements of the main results. Our main result, Theorem 2.1, is a general template for obtaining bounds (time-uniform boundary-crossing inequalities in particular) on truncation-style estimators. Corollary 2.2 then instantiates the template with particular parameters to obtain tightness for a fixed sample size. Section 3 is dedicated to the proof of Theorem 2.1. Section 4 then uses a technique known as “stitching” to extend our line-crossing inequalities to bounds which shrink to zero over time at an iterated logarithm rate. Finally, Section 5 provides several numerical experiments demonstrating the efficacy of our proposed estimator in practice.
2 Main Result
Define the mapping
[…]
Clearly, $\mathtt{Trunc}(x)x$ is just the projection of $x$ onto the unit ball in $\mathbb{B}$. Likewise, $\mathtt{Trunc}(\lambda x)x$ is the projection of $x$ onto the ball of radius $\lambda^{-1}$ in $\mathbb{B}$. We note that the truncated observations $\mathtt{Trunc}(X_{n})X_{n}$ are thus themselves bounded random variables, which are adapted to the underlying filtration $\mathcal{F}$.
As we discussed in Section 1.1, previous analyses of truncation-style estimators have relied on a bound on the raw second moment. To handle a central moment assumption, we will center our estimator around a naive mean estimate which has worse guarantees but whose effects wash out over time.
To formalize the above, our estimate of $\mu$ at time $n$ will be
$\widehat{\mu}_{n}(k)\equiv\widehat{\mu}_{n}(k,\lambda,\widehat{Z}_{k}):=\frac{1}{n-k}\sum_{k<m\leq n}\big{\{}\mathtt{Trunc}(\lambda(X_{m}-\widehat{Z}_{k}))(X_{m}-\widehat{Z}_{k})+\widehat{Z}_{k}\big{\}},$ (2.2)
where $\widehat{Z}_{k}$ is a naive mean estimate formed using the first $k$ samples and $\lambda>0$ is some fixed hyperparameter. Defining $\widehat{Z}_{0}=0$ when $k=0$, we observe that $\widehat{\mu}_{n}(0)$ is the usual truncation estimator, analyzed by Catoni and Giulini [6] in the fixed-time setting and Chugg et al. [11] in the sequential setting. To state our result, we define the constant $K_{p}$ as
$K_{p}:=\frac{1}{p/q+1}\left(\frac{p/q}{p/q+1}\right)^{p/q}\text{\leavevmode\nobreak\ where \leavevmode\nobreak\ }\frac{1}{p}+\frac{1}{q}=1,$ (2.3)
which depends on the Holder conjugate $q$ of $p$. Note that $K_{p}<1$ for all $p>1$. In fact, $\lim_{p\to 1}K_{p}=1$, $\lim_{p\to\infty}K_{p}=0$, and $K_{p}$ is decreasing in $p$. We also define the constant
$\mathfrak{C}_{p}(\mathbb{B})=\begin{cases}2^{p-1}(\frac{e^{2}-3}{4}),&\text{if}(\mathbb{B},\|\cdot\|)\text{ is a Hilbert space},\\ 2^{p+1}\left(\frac{e^{2\beta^{-1}}-2\beta^{-1}-1}{(2\beta^{-1})^{2}}\right),&\text{otherwise},\end{cases}$ (2.4)
which depends on the geometry and smoothness $\beta$ of the Banach space $(\mathbb{B},\|\cdot\|)$. We note that the function $\frac{e^{\rho}-\rho-1}{\rho^{2}}$ is strictly increasing in $\rho$, and thus that we have, for any $\beta\geq 1$:
$1.10.5<\frac{e^{2\beta^{-1}}-2\beta^{-1}-1}{(2\beta^{-1})^{2}}\leq\left(\frac{e^{2}-3}{4}\right)<$ 1.1 .
In a Hilbert space, we can exploit inner product structure to save a small multiplicative factor in defining $\mathfrak{C}_{p}(\mathbb{B})$.
Our main result is the following template for bounding the deviations of $\widehat{\mu}_{n}$ assuming some sort of concentration of $\widehat{Z}_{k}$ around $\mu$.
Theorem 2.1 (Main result).
Let $(X_{n})_{n\geq 1}$ be a sequence of random variables satisfying Assumption 1 which lie in some Banach space $(\mathbb{B},\|\cdot\|)$ satisfying Assumption 2. Suppose we use the first $k$ samples to construct $\widehat{Z}_{k}$, and suppose $\widehat{Z}_{k}$ satisfies, for any $\delta\in(0,1]$,
$\mathbb{P}(\|\mu-\widehat{Z}_{k}\|\geq r(\delta,k))\leq\delta,$ (2.5)
where $r:(0,1]\times\mathbb{N}\to\mathbb{R}_{\geq 0}$ is some function. Fix any $\delta\in(0,1]$. Decompose $\delta$ as $\delta=\delta_{1}+\delta_{2}$ where $\delta_{1},\delta_{2}>0$. Then, for any $\lambda>0$, with probability $1-\delta$, simultaneously for all $n\geq k$, we have:
$\left\|\widehat{\mu}_{n}(k,\lambda,\widehat{Z}_{k})-\mu\right\|\leq\lambda^{p-1}(\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})(v+r(\delta_{2},k)^{p})+\frac{\beta\log(2/\delta_{1})}{\lambda(n-k)}.$ (2.6)
The guarantee provided by Theorem 2.1 is a line-crossing inequality in the spirit of Howard et al. [17]. That is, if we multiply both sides by $n-k$, it provides a time-uniform guarantee on the probability that the left hand side deviation between $\widehat{\mu}_{n}$ and $\mu$ ever crosses the line parameterized by the right hand side of (2.6). If we optimize the value of $\lambda$ for a particular sample size $n^{*}$, the bound will remain valid for all sample sizes, but will be tightest at and around $n=n^{*}$. To obtain bounds that are tight for all $n$ simultaneously, one must pay an additional iterated logarithmic price in $n$. To accomplish this, Section 4 will deploy a carefully designed union bound over geometric epochs—a technique known as “stitching” [18]. However, for practical applications where the sample size is known in advance, we recommend Theorem 2.1 and its corollaries.
The assumption outlined in Equation (2.5) is that one can perform some “naive” estimation of the mean, say through a sample mean. The function $r(\delta,k)$ is a rate function that describes how quickly the naive mean estimate concentrates given $k$ samples. The rate of naive mean concentration may be slow, but as the total sample size $n$ grows, the effect of naive mean estimation will be negligible.
Next we provide a guideline on choosing $\lambda$ in Theorem 2.1 to optimize the tightness of our bound for a fixed sample size. The proof follows from manipulating the right hand side of (2.6).
Corollary 2.2.
In Theorem 2.1, consider taking
$\lambda=\left(\frac{\beta\log(2/\delta_{1})}{(\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})(n-k)(v+r^{p}(\delta_{2},k))}\right)^{1/p}.$ (2.7)
Then, with probability at least $1-\delta_{1}-\delta_{2}$, we have
$\|\widehat{\mu}_{n}(k)-\mu\|\leq 2\Big{(}(\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})(v+r^{p}(\delta_{2},k))\Big{)}^{1/p}\left(\frac{\beta\log(2/\delta_{1})}{n-k}\right)^{(p-1)/p}.$ (2.8)
In particular, as long as $k=o(n)$, $r(\delta,k)=o(1)$ and $\delta_{1},\delta_{2}=\Theta(\delta)$, we have
$\|\widehat{\mu}_{n}(k)-\mu\|=O\left(\beta v^{1/p}\left(\frac{\log(1/\delta)}{n}\right)^{(p-1)/p}\right).$ (2.9)
This is the desired rate per the discussion in Section 1.1, matching the rate of other estimators which hold under infinite variance. In particular, it matches the rates of Bubeck et al. [3] in scalar settings and Minsker [30] in Banach spaces.
We can make the asymptotic rate requirements outlined in Corollary 2.2 more concrete by looking at a couple of examples for the naive mean estimate. In particular, we can instantiate Theorem 2.1 when we take $\widehat{Z}_{k}$ to be either the sample mean or Minsker’s geometric median-of-means. The latter provides a better dependence on $\delta_{2}$ but at an additional computational cost. As we’ll see in Section 5, this benefit is apparent for small sample sizes, but washes out as $n$ grows. The details of the proof can be found in Appendix A.
Corollary 2.3.
Let $(X_{n})_{n\geq 1}$ be a sequence of random variables satisfying Assumption 1 which lie in some Banach space $(\mathbb{B},\|\cdot\|)$ satisfying Assumption 2. For some $k<n$, let $\widehat{Z}_{k}$ be the empirical mean of the first $k$ observations. Given $\delta>0$, decompose it as $\delta=\delta_{1}+\delta_{2}$ for any $\delta_{1},\delta_{2}>0$. Then, with probability $1-\delta$,
$\|\widehat{\mu}_{n}(k)-\mu\|\leq 2\beta v^{1/p}C_{p}^{1/p}\left(\frac{\log(2/\delta_{1})}{(n-k)^{1/p}}\right)^{(p-1)/p}\left(1+O\left(\frac{\beta^{p}}{\delta_{2}k^{p-1}}\right)\right)^{1/p},$ (2.10)
where $C_{p}=\mathfrak{C}_{p}(\mathbb{B})+\beta^{-1}K_{p}2^{p-1}$. If, on the other hand, $\widehat{Z}_{k}$ is the geometric median-of-means estimator with appropriate tuning parameters, then with probability $1-\delta$,
$\|\widehat{\mu}_{n}(k)-\mu\|\leq 2\beta v^{1/p}C_{p}^{1/p}\left(\frac{\log(2/\delta_{1})}{(n-k)^{1/p}}\right)^{(p-1)/p}\left(1+O\left(\frac{\beta^{p}\log(1/\delta_{2})^{p-1}}{k^{p-1}}\right)\right)^{1/p}.$ (2.11)
For instance, if $\widehat{Z}_{k}$ is taken either to be the empirical mean or a geometric median of means estimate, and if $k=k(n)=\lfloor\log_{2}(n)\rfloor$, the both terms in the big-Oh notation will approach 00 as $n\rightarrow\infty$ (that is, they will be $o(1)$). Using geometric median of means over the empirical mean provides better non-asymptotic dependence on $\delta_{2}$ (see Appendix B).
Remark 1.
We note that throughout this section we have explicitly focused on the setting $p\in(1,2]$. One may ask how the truncation-based estimator can be adapted to handle a moment bound of $\sup_{n}\mathbb{E}_{n-1}\|X_{n}\|^{p}\leq v$ for some $p>2$. Jensen’s inequality trivially yields for any $n\geq 1$
$\mathbb{E}_{n-1}\|X_{n}\|^{2}\leq\left(\mathbb{E}_{n-1}\|X_{n}\|^{p}\right)^{2/p}\leq v^{2/p},$
and thus applying Corollary 2.2 will yield a rate of $\|\widehat{\mu}_{n}(k)-\mu\|=O\left(\beta v^{1/p}\sqrt{\frac{\log(1/\delta)}{n}}\right)$.
3 Proof of Theorem 2.1
We will prove a slightly more general result that reduces to Theorem 2.1 in a special case. In this section, fix two arbitrary $\mathcal{F}$-predictable sequences, $(\widehat{Z}_{n})\in\mathbb{B}^{\mathbb{N}}$ and $(\lambda_{n})\in\mathbb{R}_{+}^{\mathbb{N}}$. Define
$Y_{m}$ $:=X_{m}-\widehat{Z}_{m},\text{ and }$ (3.1)
$\widehat{\xi}_{n}\equiv\widehat{\xi}_{n}(\lambda^{n},\widehat{Z}^{n})$ $:=\sum_{m\leq n}\lambda_{m}\big{\{}\mathtt{Trunc}(\lambda_{m}Y_{m})Y_{m}+\widehat{Z}_{m}\big{\}}.$ (3.2)
If we take $\lambda_{n}$ to be constant $\lambda$ and $\widehat{Z}_{m}=\widehat{Z}$ to be $\mathcal{F}_{0}$-measurable, then $\widehat{\xi}_{n}=n\lambda\widehat{\mu}_{n}$. However, our more general analysis allows us to consider sequences of predictable mean estimates, which will be useful in deriving the iterated logarithm corollary.
Our preliminary goal is to find, for any $\rho>0$, $(\widehat{Z}_{n})_{n\geq 1}$, and $(\lambda_{n})_{n\geq 1}$ a process $(V_{n})_{n\geq 0}\equiv(V_{n}(\lambda^{n},\rho,\widehat{Z}^{n}))_{n\geq 0}$ such that the process
$M_{n}(\lambda^{n})=\frac{1}{2}\exp\left\{\rho\bigg{\|}\widehat{\xi}_{n}-\mu\sum_{m\leq n}\lambda_{m}\bigg{\|}-V_{n}(\lambda^{n},\rho,\widehat{Z}^{n})\right\},$ (3.3)
is upper bounded by a nonnegative supermartingale. (In particular, it will be upper bounded by a nonnegative supermartingale with initial value 1, meaning that, in recent parlance, $M_{n}$ is an e-process [35].) Applying Ville’s inequality will then give us a time-uniform bound on the deviation of the process $\|(\sum_{m\leq n}\lambda_{m})^{-1}\widehat{\xi}_{n}-\mu\|$ in terms of $V_{n}(\lambda^{n})$.
To this end, we start by defining, for any $\rho>0$, the constant
$\mathfrak{C}_{p}(\mathbb{B},\rho):=\begin{cases}2^{p-1}\left(\frac{e^{2\rho}-2\rho-1}{(2\rho)^{2}}\right)&\text{if }(\mathbb{B},\|\cdot\|)\text{ is a Hilbert space},\\ 2^{p+1}\left(\frac{e^{2\rho}-2\rho-1}{(2\rho)^{2}}\right),&\text{otherwise},\end{cases}$
and note that $\mathfrak{C}_{p}(\mathbb{B})=\mathfrak{C}_{p}(\mathbb{B},\beta^{-1})$. We additionally define the variance process
$V_{n}\equiv V_{n}(\lambda^{n},\rho,\widehat{Z}^{n})=(\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B},\rho)+\rho K_{p}2^{p-1})G_{n},$
where $G_{n}\equiv G_{n}(\lambda^{n},\widehat{Z}^{n}):=\sum_{m\leq n}\lambda_{m}^{p}(v+\|\mu-\widehat{Z}_{m}\|^{p}).$ (3.4)
$(G_{n})_{n\geq 0}$ gives a weighted measure of the deviation of the naive estimates $\widehat{Z}_{1},\dots,\widehat{Z}_{n}$ from $\mu$, and $(V_{n})_{n\geq 0}$ further couples this deviation with the smoothness of the Banach space. Since it is difficult to reason about the difference between $\widehat{\xi}_{n}$ and $\mu\sum_{m\leq n}\lambda_{m}$ directly, we introduce the proxy
$\widetilde{\mu}_{n}(\lambda):=\mathbb{E}_{n-1}[\mathtt{Trunc}(\lambda Y_{n})Y_{n}]+\widehat{Z}_{n},$ (3.5)
so that $\lambda_{n}\widetilde{\mu}_{n}(\lambda_{n})$ is just the conditional mean of $n$th increment of $(\widehat{\xi}_{n})_{n\geq 0}$ given $\mathcal{F}_{n-1}$. We then separately argue about distances $\|\widehat{\xi}_{n}-\sum_{m\leq n}\lambda_{m}\widetilde{\mu}_{m}(\lambda_{m})\|$ and $\lambda_{m}\|\widetilde{\mu}_{m}(\lambda_{m})-\mu\|$. This gives us a bound on the difference $\|\widehat{\xi}_{n}-\mu\sum_{m\leq n}\lambda_{m}\|$ via the triangle inequality.
To build some intuition, if the $\lambda_{m}$ are all small, we are truncating onto a very large ball, and thus expect $\sum_{m\leq n}\lambda_{m}\|\widetilde{\mu}_{m}(\lambda_{m})-\mu\|$ to be small. However, in this case, our ability to bound $\|\widehat{\xi}_{n}-\sum_{m\leq n}\lambda_{m}\widetilde{\mu}_{m}(\lambda_{m})\|$ using concentration of measure results for bounded random variables will be weak. Likewise, if the $\lambda_{m}$ are large, $\widetilde{\mu}_{m}(\lambda_{m})$ will be far from $\mu$, but we will have greater control over the growth of $\widehat{\xi}_{n}$. Our argument in the sequel involves carefully balancing the growth of each term to give strong rates of concentration.
3.1 Step 1: Bounding $\|\widetilde{\mu}_{n}(\lambda)-\mu\|$
We need the following analytical property of $\mathtt{Trunc}$, which will be useful in bounding the truncation error with fewer than two moments. We note that the following lemma was used by Catoni and Giulini [6] for $k\geq 1$. We extend their result to hold for $k>0$.
Lemma 3.1.
For any $k>0$ and $x\in\mathbb{B}$, we have that
$1-\mathtt{Trunc}(x)\leq\frac{\|x\|^{k}}{k+1}\left(\frac{k}{k+1}\right)^{k}.$
Proof.
Fix $k>0$. It suffices to show that $f(t):=1-\frac{1\land t}{t}\leq\frac{t}{k+1}\left(\frac{k}{k+1}\right)^{k}=:g_{k}(t)$ for all $t\geq 0$. For $t\in[0,1],$ the result is obvious. For $t\geq 1$, we need to do a bit of work. First, note that $g_{k}(1)>f(1)=0$, and that both $g_{k}$ and $f$ are continuous. Further, we only have $g_{k}(t)=f(t)$ precisely when $t=\frac{k+1}{k}$. Let this value of $t$ be $t^{\ast}$. This immediately implies that $g_{k}(t)\geq f(t)$ for $t\in[1,t^{\ast}]$. To check the inequality for all $t\geq t^{\ast}$, it suffices to check that $f^{\prime}(t)<g_{k}^{\prime}(t)$. We verify this by direct computation. First, $f^{\prime}(t)=\frac{1}{t^{2}}$. Likewise, we have that $g_{k}^{\prime}(t)=t^{k-1}\left(\frac{k}{k+1}\right)^{k+1}$. Taking ratios, we see that
$\frac{g_{k}^{\prime}(t)}{f^{\prime}(t)}=t^{k+1}\left(\frac{k}{k+1}\right)^{k+1}\geq\left(\frac{k+1}{k}\right)^{k+1}\left(\frac{k}{k+1}\right)^{k+1}=1,$
proving the desired result. ∎
We can now proceed to bounding $\|\widetilde{\mu}_{n}(\lambda)-\mu\|$.
Lemma 3.2.
Let $X$ be a $\mathbb{B}$-valued random variable and suppose $\mathbb{E}_{n-1}\|X-\mu\|^{p}\leq v<\infty$. Let $\widehat{Z}_{n}$ be $\mathcal{F}_{n-1}$-predictable and $\widetilde{\mu}_{n}$ be as in (3.5). Then:
$\|\mu-\widetilde{\mu}_{n}(\lambda)\|\leq K_{p}2^{p-1}\lambda^{p-1}(v+\|\widehat{Z}_{n}-\mu\|^{p}).$
Proof.
Since $\widehat{Z}_{n}$ is predictable, we may treat it as some constant $z$ when conditioning on $\mathcal{F}_{n-1}$. Using Holder’s inequality, write
$\|\mu-\widetilde{\mu}_{n}(\lambda)\|$ $=\|\mathbb{E}_{n-1}[X_{n}]-\mathbb{E}_{n-1}[\mathtt{Trunc}(\lambda(X_{n}-z))(X_{n}-z)+z]\|$
$=\|\mathbb{E}_{n-1}[\{1-\mathtt{Trunc}(\lambda(X_{n}-z))\}(X_{n}-z)\|$
$\leq\mathbb{E}[|1-\mathtt{Trunc}(\lambda(X_{n}-z))|^{q}]^{1/q}\mathbb{E}[\|X_{n}-z\|^{p}]^{1/p},$
where $11/p+1/q=$ 1. The second expectation on the right hand side can be bounded using Minkowski’s inequality and the fact that $\|\cdot\|^{p}$ is convex for $p\geq 1$:
$\mathbb{E}_{n-1}[\|X_{n}-z\|^{p}]$ $=\mathbb{E}_{n-1}[\|X_{n}-\mu+\mu-z\|^{p}]$
$\leq 2^{p-1}\left(\mathbb{E}_{n-1}[\|X_{n}-\mu\|^{p}]+\|z-\mu\|^{p}\right)$
$\leq 2^{p-1}(v+\|z-\mu\|^{p}).$ (3.6)
Next, by Lemma 3.1, we have for any $k>0$.
$\mathbb{E}_{n-1}[|1-\mathtt{Trunc}(\lambda(X_{n}-z))|^{q}]\leq\mathbb{E}_{n-1}\left[\left(\frac{\lambda^{k}\|X_{n}-z\|^{k}}{k+1}\left(\frac{k}{k+1}\right)^{k}\right)^{q}\right].$
In particular, selecting $k=\frac{p}{q}$, we have
$\mathbb{E}_{n-1}[|1-\mathtt{Trunc}(\lambda(X_{n}-z))|^{q}]\leq K_{p}^{q}\mathbb{E}_{n-1}\left[\lambda^{p}\|X_{n}-z\|^{p}\right]\leq K_{p}^{q}\lambda^{p}2^{p-1}(v+\|z-\mu\|^{p}),$
where $K_{p}$ as defined in (2.3). Piecing everything together, we have that
$\mathbb{E}_{n-1}[|1-\mathtt{Trunc}(\lambda(X_{n}-z))|^{q}]^{1/q}\leq K_{p}\lambda^{p/q}2^{(p-1)/q}(v+\|z-\mu\|^{p})^{1/q}.$
Therefore, recalling that $p/q=p-1$, we have $\|\mu-\widetilde{\mu}_{n}(\lambda)\|\leq K_{p}\lambda^{p-1}2^{p-1}(v+\|z-\mu\|^{p})$, which is the desired result. ∎
3.2 Step 2: Bounding $\|\widehat{\xi}_{n}(\lambda^{n})-\sum_{m\leq n}\lambda_{m}\widetilde{\mu}_{m}(\lambda_{m})\|$
We can now proceed to bounding $\|\widehat{\xi}_{n}(\lambda^{n})-\sum_{m\leq n}\lambda_{m}\widetilde{\mu}_{m}(\lambda_{m})\|=\|S_{n}(\lambda^{n},\widehat{Z}^{n})\|$ where we have defined the process $S\equiv(S_{n})_{n\geq 0}$ by
$S_{n}\equiv S_{n}(\lambda^{n},\widehat{Z}^{n}):=\sum_{m=1}^{n}\lambda_{m}\left\{\mathtt{Trunc}(\lambda_{m}Y_{m})Y_{m}+\widehat{Z}_{m}-\widetilde{\mu}_{m}(\lambda_{m})\right\}.$ (3.7)
and $\widetilde{\mu}$ is as in (3.5). Note that $S$ is a martingale with respect to $\mathcal{F}$. The following proposition is the most technical result in the paper. It follows from a modification of the proof of Theorem 3.2 in Pinelis [34], combined with a Bennett-type inequality for 2-smooth separable Banach spaces presented in Pinelis [34, Theorem 3.4]. We present the full result here, even those parts found in Pinelis’ earlier work, for the sake of completeness.
Proposition 3.3.
Let $(X_{n})_{n\geq 1}$ be a process satisfying Assumption 1 and lying in a Banach space $(\mathbb{B},\|\cdot\|)$ satisfying Assumption 2. Then, the exponential process
$U_{n}\equiv U_{n}(\lambda^{n},\rho,\widehat{Z}^{n}):=\frac{1}{2}\exp\left\{\rho\left\|S_{n}(\lambda^{n},\widehat{Z}^{n})\right\|-\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B},\rho)G_{n}\right\},$
is bounded from above by a nonnegative supermartingale with initial value $1$, where $G_{n}$ is defined by (3.4).
Proof.
Fix some $n\geq 1$ and let $U_{n}=U_{n}(\lambda^{n},\widehat{Z}_{n})$. We first observe that
$\|\Delta S_{n}\|$ $=\lambda_{n}\|\mathtt{Trunc}(\lambda_{n}Y_{n})Y_{n}+\widehat{Z}_{n}-\widetilde{\mu}_{n}(\lambda_{n})\|$
$\leq\lambda_{n}\|\mathtt{Trunc}(\lambda_{n}Y_{n})Y_{n}\|+\lambda_{n}\|\widehat{Z}_{n}-\widetilde{\mu}_{n}(\lambda_{n})\|\leq 2,$
by definition of $\mathtt{Trunc}$. Let $T_{n}=\mathtt{Trunc}(\lambda_{n}Y_{n})Y_{n}$.
[…]
$\mathbb{E}_{n-1}\|\Delta S_{n}\|^{2}=\lambda_{n}^{2}\mathbb{E}_{n-1}\langle T_{n}-\mathbb{E}_{n-1}T_{n},T_{n}-\mathbb{E}_{n-1}T_{n}\rangle\leq\lambda_{n}^{2}\mathbb{E}_{n-1}\|T_{n}\|^{2}.$
Otherwise, we have
$\mathbb{E}_{n-1}\|\Delta S_{n}\|^{2}$ $\leq\lambda_{n}^{2}\{\mathbb{E}_{n-1}\left(\|T_{n}\|+\|\mathbb{E}_{n-1}T_{n}\|\right)^{2}\}$
$\leq 2\lambda_{n}^{2}\{\mathbb{E}_{n-1}\|T_{n}\|^{2}+\|\mathbb{E}_{n-1}T_{n}\|^{2}\}\leq 4\lambda_{n}^{2}\mathbb{E}_{n-1}\|T_{n}\|^{2},$
where the penultimate inequality uses that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and the final inequality follows from Jensen’s inequality. Therefore, we can write
$\mathbb{E}_{n-1}\|\Delta S_{n}\|^{2}\leq C\lambda_{n}^{2}\mathbb{E}_{n-1}\|T_{n}\|^{2},$ (3.8)
[…]
$\mathbb{E}_{n-1}\|\Delta S_{n}\|^{2}$ $\leq C\lambda_{n}^{2}\mathbb{E}_{n-1}\left[\|T_{n}\|^{p}\|T_{n}\|^{2-p}\right]$
$\leq C\lambda_{n}^{p}\mathbb{E}_{n-1}\left\|T_{n}\right\|^{p}$
$\leq C\lambda_{n}^{p}\mathbb{E}_{n-1}\left\|Y_{n}\right\|^{p}$
$\leq C\lambda_{n}^{p}2^{p-1}(v+\|\mu-\widehat{Z}_{n}\|^{p}),$ (3.9)
where the final inequality follows by the same argument used to prove (3.6) in Lemma 3.2. We have shown that the random variable $\|\Delta S_{n}\|$ is bounded and its second moment (conditioned on the past) can be controlled, which opens the door to Pinelis-style arguments (see Pinelis [34, Theorem 3.4] in particular). Define the function $\varphi:[0,1]\rightarrow\mathbb{R}_{\geq 0}$ by
$\varphi(\theta):=\mathbb{E}_{n-1}\cosh\left(\rho\|S_{n-1}+\theta\Delta S_{n}\|\right).$
In principle, the norm function need not be differentiable, and so the same applies to $\varphi$. However, Pinelis [34] proved that one may assume smoothness of the norm without loss of generality (see Pinelis [34, Remark 2.4]). Thus, a second order Taylor expansion yields
[…]
Observe that
$\varphi^{\prime\prime}(\theta)$ $\leq\rho^{2}\beta^{2}\mathbb{E}_{n-1}\left[\|\Delta S_{n}\|^{2}\cosh(\rho\|S_{n-1}\|)e^{\theta\rho\|\Delta S_{n}\|}\right]$
$\leq\rho^{2}\beta^{2}\cosh(\rho\|S_{n-1}\|)\mathbb{E}_{n-1}\left[\|\Delta S_{n}\|^{2}\right]e^{2\rho\theta},$
where the first inequality follows from the proof of Theorem 3.2 in Pinelis [34] and Theorem 3 in Pinelis [33], and the second inequality is obtained in view of $\|\Delta S_{n}\|\leq 2$.
Next, by the chain rule, we have
$\varphi^{\prime}(0)$ $=\frac{d}{d\theta}\left(\mathbb{E}_{n-1}\cosh(\rho\|S_{n-1}+\theta\Delta S_{n}\|)\right)\big{|}_{\theta=0}$
$=\mathbb{E}_{n-1}\left[\frac{d}{d\theta}\cosh(\rho\|S_{n-1}+\theta\Delta S_{n}\|)\Big{|}_{\theta=0}\right]$
$=\rho\mathbb{E}_{n-1}\left[\langle D_{f}\|f\|\big{|}_{f=S_{n-1}},\Delta S_{n}\rangle\cdot\frac{d}{dx}\cosh(x)\big{|}_{x=\|S_{n-1}\|}\right]$
$=\rho\left\langle D_{f}\|f\|\big{|}_{f=S_{n-1}},\mathbb{E}_{n-1}\Delta S_{n}\right\rangle\cdot\frac{d}{dx}\cosh(x)\big{|}_{x=\|S_{n-1}\|}$
$=0,$
where $\left\langle D_{f}\varphi(f)\mid_{f=g},y-x\right\rangle$ denotes the Gateaux derivative of $\varphi$ with respect to $f$ at $g$ in the directon of $y-x$. The final equality follows from the fact that $(S_{n})_{n\geq 0}$ is itself a martingale with respect to $(\mathcal{F}_{n})_{n\geq 1}$. Thus, leveraging that $\varphi^{\prime}(0)=0$, we have
[…]
=$\stackrel{{\scriptstyle(i)}}{{=}}\cosh(\rho\|S_{n-1}\|)\left(1+\rho^{2}\beta^{2}\left(\frac{e^{2\rho}-\rho-1}{(2\rho)^{2}}\right)\mathbb{E}_{n-1}\left[\|\Delta S_{n}\|^{2}\right]\right)$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\cosh(\rho\|S_{n-1}\|)\left(1+\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B},\rho)\lambda_{n}^{p}(v+\|\mu-\widehat{Z}_{n}\|^{p})\right)$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\cosh(\rho\|S_{n-1}\|)\exp\left\{\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B},\rho)\lambda_{n}^{p}(v+\|\mu-\widehat{Z}_{n}\|^{p})\right\},$
[…]
3.3 Step 3: Bounding $M_{n}(\lambda^{n},\rho)$
We now combine Lemma 3.2 and Proposition 3.3 to write down an explicit form for the supermartingale $M_{n}\equiv M_{n}(\lambda^{n},\rho,\widehat{Z}^{n})$ in (3.3).
Lemma 3.4.
Let $(X_{n})_{n\geq 1}$ and $(\mathbb{B},\|\cdot\|)$ be as in Proposition 3.3. Then, the process $(M_{n}(\lambda^{n},\rho,\widehat{Z}^{n}))$ defined by
$M_{n}(\lambda^{n},\rho,\widehat{Z}^{n}):=\frac{1}{2}\exp\left\{\rho\left\|\widehat{\xi}_{n}-\mu\sum_{m\leq n}\lambda_{m}\right\|-(\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B})+\rho K_{p}2^{p-1})G_{n}\right\},$
is bounded above by a nonnegative supermartingale with initial value $1$.
Proof.
Recall that $\widetilde{\mu}_{n}(\lambda)=\mathbb{E}_{n-1}[\mathtt{Trunc}(\lambda Y_{n})Y_{n}]+\widehat{Z}_{n}$. Applying the triangle inequality twice and Lemma 3.2 once, we obtain
$\left\|\widehat{\xi}_{n}-\mu\sum_{m\leq n}\lambda_{m}\right\|$ $\leq\left\|\widehat{\xi}_{n}-\sum_{m\leq n}\lambda_{m}\widetilde{\mu}_{m}(\lambda_{m})\right\|+\sum_{m\leq n}\lambda_{m}\|\widetilde{\mu}(\lambda_{m})-\mu\|$
$\leq\left\|S_{n}\right\|+K_{p}2^{p-1}\sum_{m\leq n}\lambda_{m}^{p}(v+\|\mu-\widehat{Z}_{m}\|^{p})=\left\|S_{n}\right\|+K_{p}2^{p-1}G_{n}.$
Therefore,
$M_{n}$ $=\frac{1}{2}\exp\{\rho\bigg{\|}\widehat{\xi}_{n}-\mu\sum_{m\leq n}\lambda_{m}\bigg{\|}-(\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B})+\rho K_{p}2^{p-1})G_{n}\bigg{\}}$
$\leq\frac{1}{2}\exp\left\{\rho\|S_{n}\|+\rho K_{p}2^{p-1}G_{n}-(\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B})+\rho K_{p}2^{p-1})G_{n}\right\}$
$=\frac{1}{2}\exp\left\{\rho\|S_{n}\|-\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B})G_{n}\right\},$
which is itself upper bounded by a nonnegative supermartingale with initial value 1 by Proposition 3.3. ∎
We are finally ready to prove Theorem 2.1, which follows as a consequence of the following result.
Proposition 3.5.
Let $(\mathbb{B},\|\cdot\|)$ satisfy Assumption 2 and $(X_{n}^{\prime})_{n\geq 1}$ satisfy Assumption 1 with respect to some filtration $(\mathcal{G}_{n})_{n\geq 0}$. Suppose $\widehat{Z}$ is $\mathcal{G}_{0}$-measurable and there exists some function $r:(0,1)\to\mathbb{R}_{\geq 0}$ such that, for any $\delta\in(0,1]$,
$\mathbb{P}(\|\mu-\widehat{Z}\|\geq r(\delta))\leq\delta.$ (3.10)
Fix any $\delta\in(0,1]$. Decompose $\delta$ as $\delta=\delta_{1}+\delta_{2}$ where $\delta_{1},\delta_{2}>0$. Then, for any $\rho,\lambda>0$, with probability $1-\delta$, simultaneously for all $n\geq 1$, we have:
$\left\|\widehat{\mu}_{n}-\mu\right\|\leq\lambda^{p-1}(\rho\beta^{2}\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})(v+r(\delta_{2})^{p})+\frac{\log(2/\delta_{1})}{\lambda\rho n},$ (3.11)
where $\widehat{\mu}_{n}=\frac{1}{n}\sum_{m\leq n}\{\mathtt{Trunc}(\lambda(X_{m}^{\prime}-\widehat{Z}))(X_{m}^{\prime}-\widehat{Z})+\widehat{Z}\}$.
Proof.
Let $B_{1}=\{\exists n:M_{n}(\lambda^{n})\geq 2/\delta_{1}\}$ where $(M_{n})$ is as in Lemma 3.4. By Ville’s inequality (Section 1.3), $\mathbb{P}(B_{1})\leq\delta_{1}$. Let $B_{2}=\{\|\mu-\widehat{Z}\|\geq r(\delta_{2})\}$. By assumption, $\mathbb{P}(B_{2})\leq\delta_{2}$. Set $B=B_{1}\cup B_{2}$ so that $\mathbb{P}(B)\leq\delta$. We take the sequence of predictable values $(\lambda_{n})$ in Lemma 3.4 to be constant and set $\lambda_{n}=\lambda>0$ for all $n$. On the event $B^{c}$ we have $\log(M_{n})\leq\log(2/\delta_{1})$ for all $n\geq 1$. That is, with probability $1-\delta$,
$\rho\|\widehat{\xi}_{n}-\lambda n\mu\|\leq(\rho^{2}\beta^{2}\mathfrak{C}_{p}(\mathbb{B})+\rho K_{p}2^{p-1})G_{n}+\log(2/\delta_{1}),$ (3.12)
and
$G_{n}=n\lambda^{p}(v+\|\mu-\widehat{Z}\|^{p})\leq n\lambda^{p}(v+r(\delta_{2})^{p}).$ (3.13)
Substituting (3.13) into (3.12) and dividing both sides by $n\lambda\rho$ gives the desired result. ∎
Proof of Theorem 2.1.
Given $(X_{n})_{n\geq 1}$ as in the statement of Theorem 2.1 apply Proposition 3.5 with $X_{n}^{\prime}=X_{n+k}$ and $\mathcal{G}_{n}=\mathcal{F}_{n+k}$ for all $n\geq 0$. Finally, take $\rho=\beta^{-1}$. ∎
4 Law of the Iterated Logarithm Rates
In the previous section, we derived a time-uniform, line-crossing inequality that controlled (with high probability) the deviation between a truncated mean estimator and the unknown mean. This inequality was parameterized by a scalar/truncation level $\lambda$, which, when optimized appropriately, could guarantee a width of $O(\beta v^{1/p}(\log(1/\delta)/n)^{(p-1)/p})$ with probability at least $1-\delta$ for a preselected sample size $n$. However, in many settings, one may not know a target sample size in advance and may wish to observe the data sequentially and stop adaptively at a data-dependent stopping time.
To generalize our bound to this setting, we use a technique known as stitching [18]. Stitching involves breaking “intrinsic time” (governed by the number of observed samples) into geometrically spaced epochs. In each epoch, we appropriately deploy and optimize Theorem 2.1. By carefully applying a union bound over these epochs, were obtain time-uniform rates of convergence that match the tightness of Corollary 2.2 up to a doubly-logarithmic multiplicative factor in the sample size.
In more detail, we break the number of observed samples into geometric buckets of the form $[2^{j},2^{j+1})$ for $j\geq 1$. Given a failure probability $\delta\in(0,1)$ and a “stitching function” $h:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ which satisfies $\sum_{j\geq 1}1/h(j)=1$, we allocate a total failure probability of $\delta/h(j)$ to the bound constructed in bucket $j$. Given this failure probability, in each bucket (i.e. for $n\in[2^{j},2^{j+1})$) we bound the performance of the estimator
$\widehat{\mu}_{n}(j):=\frac{1}{n-k(j)}\sum_{m=k(j)}^{n}\left\{\mathtt{Trunc}(\lambda_{j}(X_{m}-\widehat{Z}_{j}))(X_{m}-\widehat{Z}_{j})+\widehat{Z}_{j}\right\}.$
Here, $k(j)$ denotes the number of samples to use for the naive mean estimate in the $j$th epoch, $\widehat{Z}_{j}\equiv\widehat{Z}_{j}(X_{1},\dots,X_{k(j)})$ denotes the naive mean estimate using the first $k(j)$ samples, and $\lambda_{j}$ is some to-be-determined truncation level that will allow for rapid convergence. By separately bounding the deviations of the above estimators over each epoch, we can thus gauge the performance of the overall estimator
$\widehat{\mu}_{n}:=\sum_{j=1}^{\infty}\widehat{\mu}_{n}(j)\mathbbm{1}\{n\in[2^{j},2^{j+1})\}.$
We formalize this in the following theorem.
Theorem 4.1.
Let $(\mathbb{B},\|\cdot\|)$ satisfy Assumption 2 and let $(X_{n})_{n\geq 1}$ satisfy Assumption 1. Let $h:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a stitching function satisfying $\sum_{n=1}^{\infty}h(n)^{-1}\leq 1$ and let $k:\mathbb{N}_{\mathbf{0}}\rightarrow\mathbb{N}_{\mathbf{0}}$ be an arbitrary increasing function satisfying (i) $k(\lfloor\log_{2}(n)\rfloor)=o(n)$ and (ii) $k(\ell)\leq\ell/2$ for all $\ell\geq n_{0}$. Let $(\widehat{Z}_{j})_{j\geq 1}$ be given by $\widehat{Z}_{j}:=\widehat{Z}(X_{1},\dots,X_{k(j)})$, and suppose there is some rate function $r:(0,1)\times\mathbb{N}_{\mathbf{0}}\rightarrow\mathbb{R}$ satisfying
$\mathbb{P}\left(\|\widehat{Z}_{m}-\mu\|\geq r(\delta,j)\right)\leq\delta$
for any $\delta\in(0,1)$ and any $j\in\mathbb{N}_{\mathbf{0}}$.
Then, for any confidence parameter $\delta\in(0,1)$, we have with probability at least $1-\delta$,
$\|\widehat{\mu}_{n}-\mu\|\leq\mathfrak{B}_{p}(\mathbb{B})(v+r_{n}^{p})^{1/p}\left(\beta\frac{\log(4h(\lfloor\log_{2}(n)\rfloor))}{n-k(\lfloor\log_{2}(n)\rfloor)}\right)^{(p-1)/p}.$
simultaneously for all $n\geq n_{0}$, where $\widehat{\mu}_{n}$ is as defined above,
$r_{n}$ $:=r\left(h(\lfloor\log_{2}(n)\rfloor)^{-1}\delta/2,k(\lfloor\log_{2}(n)\rfloor)\right)\text{\leavevmode\nobreak\ and}$
$\mathfrak{B}_{p}(\mathbb{B})$ $:=[2^{(p-1)/p}+2^{1/p}](\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})^{1/p}.$
Before proceeding with the proof of Theorem 4.1, we show that if the naive estimate converges in the large sample limit and if sufficiently many samples are used to construct the naive estimate, then we get the same rate obtained in Corollary 2.2 up to a doubly-logarithmic factor in the sample size. The proof of the following is immediate upon noting that $(1+\lfloor\log_{2}(n)\rfloor)^{s}=O(\log_{2}(n))$ for any $s>1$.
Corollary 4.2.
In addition to the assumptions of Theorem 4.1, if we further suppose that
$r_{n}:=r\Big{(}h(\lfloor\log_{2}(n)\rfloor)^{-1}\delta,k(\lfloor\log_{2}(n)\rfloor)\Big{)}=o(1)$
and that $h$ is taken as $h(x):=(x+1)^{s}\zeta(s)$ for some $s>1$, then
$\|\widehat{\mu}_{n}-\mu\|=O\left(\beta v^{1/p}\left(\frac{\log(\log(n)/\delta)}{n}\right)^{(p-1)/p}\right).$
For instance, if we take $h$ as defined above simply set $k(j)=j$ and we take $\widehat{Z}_{j}:=j^{-1}(X_{1}+\cdots+X_{j})$ and obtain $r(\delta,k)$ using Lemma 1.1, then it is clear that we have
$r_{n}\lesssim\frac{\beta\log(n)^{1/p}v^{1/p}}{\delta\log(n)^{(p-1)/p}}=\frac{\beta v^{1/p}}{\delta\log(n)^{p-1}}\xrightarrow[n\rightarrow\infty]{}0.$
Likewise, similar convergence guarantees are provided if $\widehat{Z}_{j}$ is taken to be, say, a geometric median of medians estimator. We now prove Theorem 4.1.
Proof.
Set $\delta_{j}:=\delta/h(j)$ and set
$\mathfrak{C}_{p}(j):=\left(\beta\mathfrak{C}_{p}(\mathbb{B})+2^{p-1}K_{p}\right)\left(v+r(\delta_{j}/2,k(j)\right),$
for convenience. For any $j\geq 1$, defined the tuning parameter $\lambda_{j}$ as
$\lambda_{j}:=\left(\frac{\beta\log(4/\delta_{j})}{2^{j}\mathfrak{C}_{p}(j)}\right)^{1/p}.$
For $n\geq n_{0}$, $j\geq 1$, define the confidence bounds
$W(n,j):=\lambda_{j}^{p-1}\mathfrak{C}_{p}(j)+\frac{\beta\log(4/\delta_{j})}{\lambda_{j}(n-k(j))}$
If we define the “bad” event $B:=\left\{\exists n\geq n_{0}:\|\widehat{\mu}_{n}-\mu\|\geq W(n,j)\right\}$, it is clear from applying Theorem 2.1 that we have
$\mathbb{P}(B)$ $\leq\sum_{j\geq 1}\mathbb{P}(\exists n\in[2^{j},2^{j+1}):\|\widehat{\mu}_{n}(j)-\mu\|\geq W(n,j))$
$\leq\sum_{j\geq 1}\delta_{j}=\sum_{j\geq 1}h(j)^{-1}\delta\leq\delta.$
We now want to show the target time-uniform bound on the sequence of estimators $(\widehat{\mu}_{n})_{n\geq n_{0}}$. Operating on the “good” event $B^{c}$, letting $n\geq n_{0}$ be arbitrary, and letting $j$ denote the index satisfying $n\in[2^{j},2^{j+1})$, we have
$\|\widehat{\mu}_{n}-\mu\|$ $=\|\widehat{\mu}_{n}(j)-\mu\|$
$\leq W(n,j)=\lambda_{j}^{p-1}\mathfrak{C}_{p}(j)+\frac{\log(4/\delta_{j})}{\lambda_{j}(n-k(j))}$
$=\mathfrak{C}_{p}(j)^{1/p}\left(\beta\log(4/\delta_{j})\right)^{(p-1)/p}\left[\left(\frac{1}{2^{j}}\right)^{(p-1)/p}+\left(\frac{2^{j}}{n-k(j)}\right)^{1/p}\right]$
$=\mathfrak{C}_{p}(j)^{1/p}\left(\frac{\beta\log(4/\delta_{j})}{n-k(j)}\right)^{(p-1)/p}\left[\left(\frac{n-k(j)}{2^{j}}\right)^{(p-1)/p}+\left(\frac{2^{j}}{n-k(j)}\right)^{1/p}\right]$
$\leq\mathfrak{C}_{p}(j)^{1/p}\left(\frac{\beta\log(4/\delta_{j})}{n-k(j)}\right)^{(p-1)/p}\left[2^{(p-1)/p}+2^{1/p}\right],$
where the last inequality follows from the fact that $2^{j-1}\leq n-k(j)\leq 2^{j+1}$, since $n\geq n_{0}$ implies $k(j)\leq n/2$.
Lastly, recalling that $\delta_{j}=h(j)^{-1}\delta$ and that, by definition, $j=\lfloor\log_{2}(n)\rfloor$, we have
$\log(4/\delta_{j})\leq\log(4h(\lfloor\log_{2}(n)\rfloor)).$
Plugging everything together and expanding the definition of $\mathfrak{C}_{p}(j)$ yields that, with probability at least $1-\delta$, simultaneously for all $n\geq n_{0}$
$\|\widehat{\mu}_{n}-\mu\|$ $\leq[2^{(p-1)/p}+2^{1/p}]\left((\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})(v+r_{n}^{p})\right)^{1/p}\left(\frac{\beta\log(4h(\lfloor\log_{2}(n)\rfloor))}{n-k(j)}\right)^{(p-1)/p}$
$=\mathfrak{B}_{p}(\mathbb{B})\left(\frac{\beta\log(4h(\lfloor\log_{2}(n)\rfloor))}{n-k(\lfloor\log_{2}(n)\rfloor)}\right)^{(p-1)/p},$
thus proving the claimed result. ∎
5 Bound Comparison and Simulations
In the above sections, we argued that the truncated mean estimator, when appropriately optimized, could obtain a distance from the true mean of $O\big{(}v^{1/p}\left(\log(1/\delta)/n\right)^{(p-1)/p}\big{)}$ with high probability. In particular, this rate matched that of the geometric median-of-means estimator due to Minsker [30]. In this section, we study the empirical instead of theoretical performance of our bounds and estimator.
Comparing Tightness of Bounds
In Figure 1, we compare the confidence bounds obtained for our truncation-based estimators optimized for a fix sample size (Corollary 2.3) against other bounds in the literature. Namely, we compare against geometric median-of-means [30], the sample mean, and (in the case a shared covariance matrix exists for observations) the tournament median-of-means estimator [26]. We plot the natural logarithm of the bounds against the logarithm base ten of the sample sizes $n$ for $n\in[10^{2},10^{10}]$ and for $p\in\{1.25,1.5,1.75,2.0\}$. We assume $\delta=10^{-4}$ and $v=1$. For truncation-based estimates, we assume $k=\lfloor n/10\rfloor$ samples are used to produce the initial mean estimate and the remaining $n-k$ are used for the final mean estimate. We plot the resulting bounds for when the initial mean estimate is either computed using the sample mean or geometric median-of-means. For the tournament median-of-means estimate, we assume observations take their values in $\mathbb{R}^{d}$ for $d=100$, and that the corresponding covariance matrix is the identity $\Sigma=I_{d}/d$.
[…]
(a) $p=2.0$
(b) $p=1.75$
(c) $p=1.5$
(d) $p=1.25$
Figure 1: For $p\in\{1.25,1.5,1.75\}$, we plot the tightness of optimized bounds associated with the sample mean, geometric median-of-means (Geo-MoM), truncation with initial sample mean estimate, and truncation with initial Geo-MoM estimate. We assume $n\in[10^{2},10^{10}]$, $v=1.0$, $\delta=10^{-4}$, and $k=n/10$. In the case $p=2.0$, we assume a shared covariance matrix $\Sigma$ exists so we can plot the tournament median-of-means bounds assuming $\lambda_{\max}(\Sigma)=v/d$ and $d=100$.
Performance of Estimators on Simulated Data
[…]
To compute the number of folds for geometric median-of-means, we follow the parameter settings outlined in Minsker [30] and assume a failure probability of $\delta=10^{-4}$ (although we are not constructing confidence intervals, the failure probability guides how to optimize the estimator). See Appendix B for further discussion on this estimator. Once again, we consider the truncated mean estimator centered at both the sample mean and a geometric median-of-means estimate. We always use $k=\lfloor\sqrt{n}\rfloor$ samples to construct the initial estimate, and produce a plot for hyperparameter $\lambda\in[0.0005,0.005,0.05,0.5]$.
(a) $\lambda=0.0005$
(b) $\lambda=0.005$
(c) $\lambda=0.05$
(d) $\lambda=0.5$
Figure 2: We compare the empirical distributions of distance between the mean estimate and the true mean for a variety of estimators. We generate $n=10^{6}$ i.i.d. samples in $\mathbb{R}^{10}$ as outlined above, and use $k=\lfloor\sqrt{10^{6}}\rfloor$ samples to construct initial mean estimates. We compute these estimates of 250 runs. For truncation-based estimates, we consider $\lambda\in[0.0005,0.005,0.05,0.5]$. We only include the sample mean in the first plot for readability.
We construct these estimators over 250 independent runs and then construct box and whisker plots summarizing the empirical distance between the estimators and the true mean. The boxes have as a lower bound the first quartile $Q1$, in the middle the sample median $M$, and at the top the third quartile $Q3$. The whiskers of the plot are given by the largest and smallest point falling within $M\pm 1.5\times(Q3-Q2)$, respectively. All other points are displayed as outliers. We only include the sample mean in the first plot as to not compress the empirical distributions associated with other estimates.
As expected, the sample mean suffers heavily from outliers. For $\lambda\in\{0.0005,0.005\}$ (corresponding to truncation at large radii), the geometric median-of-means estimate is roughly two times closer to the mean than either truncation-based estimate. In the setting of aggressive truncation ($\lambda\in\{0.05,0.5\}$), the truncated mean estimator centered at the geometric median-of-means initial estimate offers a significantly smaller distance to the true mean than just geometric median-of-means alone. The truncated estimate centered at the sample mean performs similarly for $\lambda=0.05$, but suffers heavily from outliers when $\lambda=0.5$. Interestingly, the recommended truncation level for optimizing tightness at $n=100,000$ samples is $\lambda\approx 0.0004$ per Corollary 2.2. Our experiments reflect that one may want to truncate more aggressively than is recommended in the corollary. In practice, one could likely choose an appropriate truncation level through cross-validation.
6 Summary and Conclusion
In this work, we presented a novel analysis of a simple truncation/threshold-based estimator of a heavy-tailed mean in smooth Banach spaces, strengthening the guarantees on such estimators that currently exist in the literature. In particular, we allow for martingale dependence between observations, replace the assumption of finite variance with a finite $p$-th moment for $1<p\leq 2$, and let the centered $p$-th moment be bounded instead of the raw $p$-th moment (thus making the estimator translation invariant). Our bounds are also time-uniform, meaning they hold simultaneously for all sample sizes. We provide both a line-crossing inequality that can be optimized for a particular sample size (but remains valid at all times), and a bound whose width shrinks to zero at an iterated logarithm rate. Experimentally, our estimator performs quite well compared to more computationally intensive methods such as geometric median-of-means, making it an appealing choice for practical problems.
There are several important open directions related to the work in this paper. We note that to ensure our estimator obtained a rate of $O\left(\beta v^{1/p}(\log(\delta^{-1})/n)^{(p-1)/p}\right)$, we needed to select the truncation level $\lambda$ to explicitly depend on the moment bound $v$. This is in contrast to point estimators such as geometric median of means, whose parameter need not make explicit use of moment bounds to obtain the same rate. Thus, it would be interesting to see if one could develop point estimator analogue of the estimator considered in this paper, or perhaps an estimator that could automatically adapt to an unknown $p$th moment bound $v$.
Further, in the setting of finite-dimensional Euclidean space where $p=2$ and the data are i.i.d. with shared covariance matrix $\Sigma$, estimators have been constructed (see Lugosi and Mendelson [26, 24]) obtaining a high-probability rate of
[…]
$\|\widehat{\mu}_{n}-\mu\|\lesssim\beta v^{1/p}\left(\frac{1}{n}\right)^{(p-1)/p}+\beta w^{1/p}\left(\frac{\log(1/\delta)}{n}\right)^{(p-1)/p}$
for some appropriately defined estimator $\widehat{\mu}_{n}$ of $\mu$. We leave obtaining such a bound as interesting future work.
Acknowledgments
DMT acknowledges the support of a fellowship from ‘la Caixa’ Foundation (ID 100010434), code number LCF/BQ/EU22/11930075. BC and AR acknowledge support from NSF grant IIS-2229881. BC was supported in part by the NSERC PGS D program, grant no. 567944.
Appendix A Omitted Proofs
Definition A.1.
Let $(X_{n})_{n\geq 1}$ be a sequence of $\mathbb{B}$-valued random variables on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and let $(\mathcal{F}_{n})_{n\geq 0}$ denote its natural filtration. We say $(X_{n}^{\prime})_{n\geq 1}$ is a decoupled tangent process for $(X_{n})_{n\geq 1}$ if the following hold.
  1. 1.
    $X_{n}=_{d}X_{n}^{\prime}\mid\mathcal{F}_{n-1}$ for all $n\geq 1$, and
  2. 2.
    For any $n\geq 1$, $X_{1}^{\prime},\dots X_{n}^{\prime}$ are conditionally independent given $\mathcal{F}_{\infty}:=\bigcup_{n=0}^{\infty}\mathcal{F}_{n}$.
Proposition A.2 (Marcinkiewicz-Zygmund/von Bahr-Esseen in Smooth Banach Spaces).
Let $(X_{n})_{n\geq 1}$ be random variables in a $\beta$-smooth Banach space such that $\mathbb{E}_{n-1}X_{n}=0$ and $\mathbb{E}\|X_{n}\|^{p}<\infty$ for all $n\geq 0$ and some $p\in(1,2]$. Letting $(S_{n})_{n\geq 0}$ be $S_{n}:=X_{1}+\cdots+X_{n}$, it follows that:
  1. 1.
    If the $(X_{n})_{n\geq 1}$ are independent, we have
    $\mathbb{E}\|S_{n}\|^{p}\leq 2^{p}\beta^{p}\mathbb{E}\left[\left(\sum_{m=1}^{n}\|X_{m}\|^{2}\right)^{p/2}\right]\leq 2^{p}\beta^{p}\sum_{m=1}^{n}\mathbb{E}\|X_{m}\|^{p}.$
  2. 2.
    Otherwise,
    $\mathbb{E}\|S_{n}\|^{p}\leq 2^{p}\beta^{p}\sum_{m=1}^{n}\mathbb{E}\|X_{m}\|^{p}.$
Proof.
Let us first assume that $(X_{n})_{n\geq 1}$ are independent. In such a case, the result follows from combining symmetrization arguments and Jensen’s inequality. More specifically, let $(\epsilon_{n})_{n\geq 1}$ be a sequence of i.i.d. Rademacher random variables that are also independent of $(X_{n})_{n\geq 1}$. [^3] A random variable $\epsilon$ follows a Rademacher distribution if $\mathbb{P}(\epsilon=1)=\mathbb{P}(\epsilon=-1)=\frac{1}{2}$. It suffices to observe that
$\mathbb{E}\left\|\sum_{m=1}^{n}X_{m}\right\|^{p}$ $=\mathbb{E}\left\|\sum_{m=1}^{n}(X_{n}-\mathbb{E}X_{n})\right\|^{p}$
$\leq\mathbb{E}\left\|\sum_{m=1}^{n}(X_{m}-X_{m}^{\prime})\right\|^{p}$ Jensen’s inequality
$=\mathbb{E}\left\|\sum_{m=1}^{n}\epsilon_{m}(X_{m}-X_{m}^{\prime})\right\|^{p}$ Symmetry
$\leq 2^{p}\mathbb{E}\left\|\sum_{m=1}^{n}\epsilon_{m}X_{m}\right\|^{p}$ $\text{Convexity of }\|\cdot\|^{p}$
$\leq 2^{p}\beta^{p}\mathbb{E}\left(\sum_{m=1}^{n}\|X_{m}\|^{2}\right)^{p/2},$
to prove the first inequality. The second inequality follows because $(a+b)^{q}\leq a^{q}+b^{q}$ for $a,b\geq 0,q\in(0,1]$.
If $(X_{n})_{n\geq 1}$ are not independent, let $(X_{n}^{\prime})_{n\geq 1}$ be a decoupled tangent sequence for $(X_{n})_{n\geq 1}$ (per Definition A.1). Let $(\mathcal{G}_{n})_{n\geq 0}$ denote the natural filtration associated with $(X_{n}^{\prime})_{n\geq 1}$ and let $(\mathcal{H}_{n})_{n\geq 0}$ denote the join of $\mathcal{G}$ and $\mathcal{F}$. Further, let $(\epsilon_{n})_{n\geq 0}$ be a sequence of i.i.d. Rademacher random variables independent of $\mathcal{H}_{\infty}$. First, by construction, we note that $X_{n}-X_{n}^{\prime}$ is conditionally symmetric given $\mathcal{F}_{n-1}$. Thus, we have
$\epsilon_{n}(X_{n}-X_{n}^{\prime})=_{d}X_{n}-X_{n}^{\prime}\mid\mathcal{F}_{n-1}.$
We prove the desired claim inductively. For $n=1$, the result is trivially true. Next, observe that we have:
$\mathbb{E}\|S_{n}\|^{p}$ $=\mathbb{E}\|S_{n-1}+X_{n}\|^{p}$
$=\mathbb{E}\|S_{n-1}+X_{n}-\mathbb{E}(X_{n}^{\prime}\mid\mathcal{F}_{n-1})\|^{p}$ (Definition of $X_{n}^{\prime}$)
$\leq\mathbb{E}\left[\mathbb{E}\left(\|S_{n-1}+X_{n}-X_{n}^{\prime}\|^{p}\mid\mathcal{F}_{n-1}\right)\right]$ (Jensen’s inequality)
$=\mathbb{E}\left[\mathbb{E}\left(\|S_{n-1}+\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{p}\mid\mathcal{F}_{n-1}\right)\right]$ (Conditional Symmetry)
$=\mathbb{E}\left[\|S_{n-1}+\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{p}\right]$ (Tower Rule)
$=\mathbb{E}\left[\mathbb{E}\left(\|S_{n-1}+\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{p}\mid\mathcal{H}_{n}\right)\right]$ (Tower Rule)
$\leq\mathbb{E}\left[\mathbb{E}\left(\|S_{n-1}+\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{2}\mid\mathcal{H}_{n}\right)^{p/2}\right]$ (Jensen’s inequality)
=:()=:($\star$).= : ( ⋆ ) .
Next, define the function $g(\theta):=\mathbb{E}\left(\|S_{n-1}+\theta\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{2}\mid\mathcal{H}_{n}\right)$. In principle, the norm function need not be differentiable, and the same applies to $g$. Nonetheless, and analogously to the proof of Proposition 3.3, one can invoke Pinelis [34, Remark 2.4] to assume smoothness of the norm without loss of generality, alongside the properties exhibited in Pinelis [33, Lemma 2.2]. It follows that
$g^{\prime}(0)$ $=\mathbb{E}\left(\frac{d}{d\theta}\|S_{n-1}+\theta\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{2}\mid\mathcal{H}_{n}\right)$
$=2\mathbb{E}\left(\|S_{n-1}\|\langle D_{v}\|v\||_{v=S_{n-1}},\epsilon_{n}(X_{n}-X_{n}^{\prime})\rangle\mid\mathcal{H}_{n}\right)$
$=2\|S_{n-1}\|\left\langle D_{v}\|v\|_{v=S_{n-1}},\mathbb{E}[\epsilon_{n}(X_{n}-X_{n}^{\prime})\mid\mathcal{H}_{n}]\right\rangle$
$=0.$
Likewise, from the second property exhibited in Pinelis [33, Equation 2.4, Lemma 2.2], one has
$g^{\prime\prime}(\theta)\leq 2\beta^{2}\mathbb{E}(\|\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{2}\mid\mathcal{H}_{n}).$
Thus, putting these ideas together, we have
[…]
$\leq\mathbb{E}\left(\|S_{n-1}\|^{2}\mid\mathcal{H}_{n}\right)+\beta^{2}\mathbb{E}\left(\|\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{2}\mid\mathcal{H}_{n}\right)$
$=\|S_{n-1}\|^{2}+\beta^{2}\|X_{n}-X_{n}^{\prime}\|^{2}.$
Plugging this into ()($\star$)( ⋆ ) above, we obtain
(($\star$)( ⋆ ) $=\mathbb{E}\left[\mathbb{E}\left(\|S_{n-1}+\epsilon_{n}(X_{n}-X_{n}^{\prime})\|^{2}\mid\mathcal{H}_{n}\right)^{p/2}\right]$
$\leq\mathbb{E}\left(\|S_{n-1}\|^{2}+\beta^{2}\|X_{n}-X_{n}^{\prime}\|^{2}\right)^{p/2}$
$\leq\mathbb{E}\|S_{n-1}\|^{p}+\beta^{p}\mathbb{E}\|X_{n}-X_{n}^{\prime}\|^{p}$ $(a+b)^{q}\leq a^{q}+b^{q}\;\forall a,b\geq 0,q\in(0,1]$
[…]
concluding the proof. ∎
Proof of Lemma 1.1.
Proposition A.2 provides a bound on the $p$-th moment of sums in smooth Banach spaces. The proof of Lemma 1.1 now follows as a corollary. Applying Markov’s/Chebyshev’s inequality yields that, for any $t>0$, we have
$\mathbb{P}(\|\widehat{\mu}_{n}-\mu\|\geq t)\leq\frac{\mathbb{E}\|\widehat{\mu}_{n}-\mu\|^{p}}{t^{p}}=\frac{\mathbb{E}\left\|\sum_{m=1}^{n}(X_{m}-\mu)\right\|^{p}}{n^{p}t^{p}}.$
By Proposition A.2, we see that
$\mathbb{E}\left\|\sum_{m=1}^{n}(X_{m}-\mu)\right\|^{p}$ $\leq 2^{p}\beta^{p}\sum_{m=1}^{n}\mathbb{E}\|X_{m}\|^{p}\leq n2^{p}\beta^{p}v,$
where the final inequality follows by Assumption 1. Thus, we have
$\mathbb{P}(\|\widehat{\mu}_{n}-\mu\|\geq t)\leq\frac{n^{1-p}v2^{p}\beta^{p}}{t^{p}}.$
To meet the target failure probability $\delta$, we simply take $t=\delta^{-1/p}v^{1/p}2\beta n^{-(p-1)/p}$, which proves the desired result. ∎
Proof of Corollary 2.3.
Let $\widehat{Z}_{k}$ be the empirical mean on the first $k-1$ observations. By Lemma 1.1, we can take $r^{p}(\delta,k)=\frac{v2^{p}\beta^{p}}{\delta(k-1)^{p-1}}$. Taking $\lambda$ as in Corollary 2.2 gives that with probability $1-\delta$, for all $n>k$,
$\|\widehat{\mu}_{n}(k)-\mu\|$ $\leq 2\Big{(}(\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})(v+r^{p}(\delta_{2},k))\Big{)}^{1/p}\left(\frac{\beta\log(2/\delta_{1})}{n-k}\right)^{(p-1)/p}$
$=2\beta C_{p}^{1/p}(v+r^{p}(\delta_{2},k))^{1/p}\left(\frac{\log(2/\delta_{1})}{n-k}\right)^{(p-1)/p}$
$=2\beta C_{p}^{1/p}\left(v+\frac{v2^{p}\beta^{p}}{\delta k^{p-1}}\right)^{1/p}\left(\frac{\log(2/\delta_{1})}{n-k}\right)^{(p-1)/p}$
$=2\beta(vC_{p})^{1/p}\left(1+O\left(\frac{\beta^{p}}{\delta_{2}k^{p-1}}\right)\right)^{1/p}\left(\frac{\log(2/\delta_{1})}{n-k}\right)^{(p-1)/p}$
where $C_{p}=\mathfrak{C}_{p}(\mathbb{B})+\beta^{-1}K_{p}2^{p-1}$. When $\widehat{\mu}$ the geometric median-of-means estimator, Appendix B shows that we may take
$r^{p}(\delta_{2},k)=\frac{11^{p}\beta^{p}}{0.1}v\left(\frac{3.5\log(1/\delta)+1}{k}\right)^{p-1}.$
Again, taking $\lambda$ as in Corollary 2.2 gives that with probability $1-\delta$, for all $n>k$,
$\|\widehat{\mu}_{n}(k)-\mu\|$ $\leq 2\Big{(}(\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})(v+r^{p}(\delta_{2},k))\Big{)}^{1/p}\left(\frac{\beta\log(2/\delta_{1})}{n-k}\right)^{(p-1)/p}$
$\leq 2\beta(vC_{p})^{1/p}\left(1+O\left(\frac{\beta^{p}\log(1/\delta)^{p-1}}{k^{p-1}}\right)\right)^{1/p}\left(\frac{\log(2/\delta_{1})}{n-k}\right)^{(p-1)/p},$
as desired. ∎
Appendix B Geometric Median-of-Means in Banach Spaces
While Minsker [30] studied mean estimation in smooth Banach spaces, his examples weren’t stated explicitly for Banach spaces nor for the case of infinite variance. Here we show that his geometric median-of-means estimator, when paired with empirical mean, achieves rate $O(v^{1/p}(\log(1/\delta)/n)^{\frac{p-1}{p}})$, the same as our rate. As usual, we assume we are working in a $\beta$-smooth Banach space.
Minsker [30, Theorem 3.1] provides the following bound. Let $\widehat{\mu}_{1},\dots,\widehat{\mu}_{B}$ be a collection of independent estimators of the mean $\mu$. Fix $\alpha\in(0,1/2)$. Let $0<\gamma<\alpha$ and $\epsilon>0$ be such that for all $b$, $1\leq b\leq B$, we have
$\mathbb{P}(\|\mu-\widehat{\mu}_{b}\|>\epsilon)\leq\gamma.$ (B.1)
Let $\widehat{\mu}=\text{median}(\widehat{\mu}_{1},\dots,\widehat{\mu}_{B})$ be the geometric median, defined as
$\widehat{\mu}:=\arg\min_{\mu\in\mathbb{B}}\sum_{b=1}^{B}\|\widehat{\mu}_{b}-\mu\|.$
Then
$\mathbb{P}(\|\mu-\widehat{\mu}\|>C_{\alpha}\epsilon)\leq\exp(-B\psi(\alpha;\gamma)),$ (B.2)
where
$\psi(\alpha;\gamma)=(1-\alpha)\log\frac{1-\alpha}{1-\gamma}+\alpha\log\frac{\alpha}{\gamma},$
and
$C_{\alpha}=\frac{2(1-\alpha)}{1-2\alpha}.$
We will optimize Minsker’s bound by taking the same optimization parameters as in his paper. That is, we will set $\alpha_{\ast}:=\frac{7}{18}$ and $\gamma_{\ast}:=0.1$ and will set the number of naive mean estimators $B$ to be given by
$B:=\left\lfloor\frac{\log(1/\delta)}{\psi(\alpha_{\ast};\gamma_{\ast})}\right\rfloor+1\leq 3.5\log(1/\delta)+1,$
which provides an overall failure probability of at most $\exp(-B\psi(\alpha_{\ast};\gamma_{\ast}))\leq\delta.$ Lemma 1.1 gives
$\mathbb{P}(\|\widehat{\mu}_{b}-\mu\|\geq\epsilon)\leq\frac{v2^{p}\beta^{p}}{\epsilon^{p}(n/B)^{p-1}}\leq\gamma_{\ast},$
if we take
$\epsilon=\left(\frac{B^{p-1}v2^{p}\beta^{p}}{\gamma_{*}n^{p-1}}\right)^{1/p}.$
Therefore, we obtain that with probability $1-\delta$,
$\|\widehat{\mu}-\mu\|$ $\leq C_{\alpha_{\ast}}\epsilon=\frac{4\beta(1-\alpha_{\ast})}{1-2\alpha_{\ast}}\left(\frac{B^{p-1}}{\gamma_{\ast}n^{p-1}}\right)^{1/p}v^{1/p}\leq\frac{11\beta}{0.1^{1/p}}v^{1/p}\left(\frac{3.5\log(1/\delta)+1}{n}\right)^{1-1/p}.$
Appendix C Noncentral moment bounds
For completeness, we state our bound when we assume only a bound on the raw (uncentered) $p$-th moment of the observations. This was the setting studied by Catoni and Giulini [6]. We replace assumption 1 with the following:
Assumption 3.
We assume $(X_{n})_{n\geq 1}$ are a sequence of $\mathbb{B}$-valued random variables adapted to a filtration $(\mathcal{F}_{n})_{n\geq 0}$ such that
  1. (1)
    $\mathbb{E}(X_{n}\mid\mathcal{F}_{n-1})=\mu$, for all $n\geq 1$ and some unknown $\mu\in\mathbb{B}$, and
  2. (2)
    $\sup_{n\geq 1}\mathbb{E}\left(\|X_{n}\|^{p}\mid\mathcal{F}_{n-1}\right)\leq v<\infty$ for some known $p\in(1,2]$ and some known constant $v>0$.
With only the raw moment assumption, we do not try and center our estimator. Instead we deploy $\widehat{\mu}_{n}(0,\lambda,0)=\frac{1}{n}\sum_{m\leq n}\mathtt{Trunc}(\lambda X_{m})X_{m}$. With this estimator we obtain the following result, which achieves the same rate as Catoni and Giulini [6] and Chugg et al. [11].
Theorem C.1.
Let $X_{1},X_{2},\dots$ be random variables satisfying Assumption 3 which live in some Banach space $(\mathbb{B},\|\cdot\|)$ satisfying Assumption 2. Fix any $\delta\in(0,1]$. Then, for any $\lambda>0$, with probability $1-\delta$, simultaneously for all $n\geq 1$, we have:
$\left\|\widehat{\mu}_{n}(0,\lambda,0)-\mu\right\|\leq 2v\lambda^{p-1}(\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})+\frac{\beta\log(2/\delta)}{\lambda n}.$ (C.1)
Moreover, if we want to optimize the bound at a particular sample size $n$ and we set $\rho=\beta^{-1}$
$\lambda=\left(\frac{\log(2/\delta)\beta}{2nv(\beta\mathfrak{C}_{p}(\mathbb{B})+K_{p}2^{p-1})}\right)^{1/p},$
then with probability $1-\delta$,
$\left\|\widehat{\mu}_{n}(0,\lambda,0)-\mu\right\|\leq 2(2v(\beta\mathfrak{C}_{p}(B)+K_{p}2^{p-1}))^{1/p}\left(\beta\frac{\log(2/\delta)}{n}\right)^{(p-1)/p}.$ (C.2)
Proof.
Apply Theorem 2.1 with $k=0$ and $\widehat{Z}_{k}=0$. Then note that we can take $r(\delta,0)=v^{1/p}$ for all $\delta$ since $\|\mu\|\leq(\mathbb{E}\|X\|^{p})^{1/p}\leq v^{1/p}$ by Jensen’s inequality. ∎